OpenStudy (aamey):

Find an angle between 0 and 2pi that is coterminal with the given angles. a) 17pi/ 5 b)-7pi/3 c)55pi/2 d)11pi/7

OpenStudy (anonymous):

For positive angles, you want to subtract \(2\pi\) until you get a negative angle, and the smallest non negative angle is the answer.

OpenStudy (anonymous):

For negative angles you add \(2\pi\) and your first non negative angle is the result.

OpenStudy (aamey):

@wio I did that for 17pi/5 and got a negative answer but i don't know how to write down the answer correctly

OpenStudy (aamey):

@wio oh jk I got it!

OpenStudy (anonymous):

\[ \frac{17\pi}{5} - 2\pi = \frac{17\pi}{5} -\frac{10\pi}{5} = \frac{7\pi}{5} \]

OpenStudy (anonymous):

So this angle is less than \(2\pi\) and greater than \(0\).